• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 20:39:21 on March 27, 2012

    Entry number 367366

    keyword=Right septum limit and tonights optics plan

    The estimated right septum current for carbon elastic at (2.251 GeV/c) is 806Amps. We got approval from Jack to go up to 810A, no more. This will limit us tonight to half of our original optics plan (which include delta scan). Jixie will do whatever possible with 810A and then move on to production.

    Tomorrow we will figure out how to get the higher current runs done for the optics.

    A copy of this log entry has been emailed to: jpchen,slifer,jixie,min